# Learning How to Vote With Principles Axiomatic Insights Into the Collective Decisions of Neural Networks

### v1

@@ 3.2 Axioms
We define axioms as functions that map a voting rule and a preference profile to a value in $\{-1,1,0\}$, where 00 means that the axiom is not applicable, $-1$ means that the desideratum is violated, and $1$ that it is satisfied. The satisfaction degree of an axiom is the ratio of the number of sampled profiles in which the axiom is satisfied to the number of sampled profiles in which it is applicable. We focus on axioms that capture basic and diverse normative properties of a voting rule $F$.
@@
All voting rules defined above satisfy anonymity and neutrality, as well as the Pareto principle, for all preference profiles. They all violate independence for some preference profile. Copeland, Split Cycle, Stable voting, Blacks rule and Weak Nanson always satisfy the Condorcet principle.
3.3 Distributions of Preference Profiles
Specifying the distribution of preference data is essential to studying the voting behavior of a society. To ensure that our results are independent of the specific choice of the distribution, we employ four different ones. (For a detailed comparison of the various voting distributions, see Boehmer et al., (2024).)
@@ 3.3 Distributions of Preference Profiles
Impartial Culture (IC) assumes that all preference profiles have the same probability of appearing. Each preference of a voter in a profile is sampled uniformly at random. The Mallows distribution (Mallows,, 1957) fixes a reference ranking $P$ and assumes that each voter’s preference is close to that ranking. Closeness to the reference ranking is defined using the Kendall-tau distance, parameterized by a dispersion parameter $\phi\in(0,1]$. This distribution reduces to IC when $\phi=1$ and concentrates all mass on $P$ as $\phi$ tends to 0.
@@
The IC and Mallows distributions are complementary: IC is simplistic and widely employed in theoretical works on voting rules; it captures an extreme case with no correlation between preferences of voters. Mallows is often employed in numerical studies of voting rules that use artificial data but wish to capture more realistic voting scenarios (Caragiannis and Micha,, 2017; Lee et al.,, 2014).
@@
The next two distributions also capture more intricate relationships between the preferences in a profile. According to the 2D-Euclidean distribution, voters and alternatives are distributed randomly in 2-dimensional Euclidean space, and the closer an alternative is to a voter the more the voter prefers that alternative. Finally, the Urn distribution (Eggenberger and Pólya,, 1923) generates a profile given a parameter $\alpha\in[0,\infty)$. Voters randomly draw their ranking from an urn. Initially, the urn includes all possible rankings over the alternatives. After a voter randomly draws from the urn, we add to the urn $\alpha n!$ copies of that ranking. When $\alpha=0$, this reduces to IC.
4 Method
@@
To answer our research questions, we develop the axiomatic deep voting framework, visualized in Figure 3. It is built around a neural network, which is a function $f_{w}:\mathbb{R}^{i}\to\mathbb{R}^{j}$ parametrized by wights $w\in\mathbb{R}^{k}$. We will instantiate this with three different neural network architectures (see Section 4.1). Every profile $\boldsymbol{P}$ is mapped, via an encoding function $e$ (see Section 4.2), to a vector $x=e(\boldsymbol{P})\in\mathbb{R}^{i}$, for which the neural network produces an output $\hat{y}\in\mathbb{R}^{j}$. [^3] For our third architecture, the encoding function is part of the neural network, while for the first two it is independent (see Section 4.2); hence we treat $e$ as a separate entity here. The decoding function $d$ (see Section 4.3) turns this output into a winning set $S=d(\hat{y})$. Thus, this setup realizes the voting rule:
@@
Figure 3: The axiomatic deep voting architecture.
4.1 Architectures
@@
First, multi-layer perceptrons (MLPs)—also known as feed-forward neural network—are the classic deep neural net (see, e.g., Goodfellow et al.,, 2016, ch. 6). They consist of an input layer of neurons, one or more hidden layers, and an output layer.
Second, convolutional neural networks (CNNs) are a standard architecture to process grid-like input data such as images (see, e.g., Goodfellow et al.,, 2016, ch. 9), and in our case profiles. Compared to MLPs, they additionally use so-called convolutional layers to to capture local, invariant patterns in the input.
Third, we device an architecture that satisfies the anonymity axiom by design: We view profiles as sentences whose words are the rankings. We use the word embedding algorithm Word2vec (Mikolov et al.,, 2013) to map each ranking to a high-dimensional embedding vector. These vectors are averaged—hence we get anonymity—and an MLP then classifies this average into a winning set. This combined architecture we call here word embedding classifiers (WECs).
4.2 Encoding
@@
How should we encode $\tilde{\boldsymbol{P}}$ so it can be inputted to a neural network? The most straightforward way is to read each alternative $a_{r}^{s}\in M$ as the number that it is and the padding symbol $\sim$ as, say, $-1$. Then the matrix $\tilde{\boldsymbol{P}}$ is regarded as a vector of dimension $m_{\mathrm{max}}n_{\mathrm{max}}$. However, this does not perform well, so, following Anil and Bao, (2021), we represent an alternative not as a number but as a one-hot vector. For $a\in\{0,\ldots,m_{\mathrm{max}}-1\}$, let $\overline{a}$ be the vector of length $m_{\mathrm{max}}$ that is $1$ at position $a$ and 00 everywhere else. For the padding symbol, let $\overline{\sim}$ be the vector of length $m_{\mathrm{max}}$ that is 00 everywhere. We write $\overline{\boldsymbol{P}}=(\overline{\tilde{a}_{r}^{s}})_{r,s}$.
@@
The encoding function for WECs regards the profile $\boldsymbol{P}=(P_{1},\ldots,P_{n})$ as a sentence with words $P_{i}$. We train it to embed these words into vectors of a fixed high dimension. Thus, unlike the previous encoding functions, this one is not separate from the neural network but rather forms the first layer of the WEC, with the remaining layers processing the embedding vectors. More precisely, we first pre-train the embeddings as follows. For a given corpus size $c$, we sample $c$-many profiles from a given distribution of profiles (e.g., IC) to form our corpus (i.e., a set of sentences). The rankings occurring in the profiles form the vocabulary of this corpus, to which we add the unk token (to later represent unknown rankings, i.e., rankings that are not in the vocabulary) and the pad token (to pad a profile to length $n_{\mathrm{max}}$). Using Word2vec, we train embeddings which represent words in the vocabulary as vectors. When instantiating the WEC architecture, these embeddings form the first layer: it maps the profile $(P_{1},\ldots,P_{n})$ to the corresponding embedding vectors $(v_{1},\ldots,v_{n})$. The next layer averages these vectors into a single vector $v$, followed by several linear layers ending with the output layer.
4.3 Decoding
Given a profile $\boldsymbol{P}$ as input, all neural network architectures produce as output the logits $\hat{y}=(\hat{y}_{0},\ldots,\hat{y}_{m_{\mathrm{max}}})$ in $\mathbb{R}^{m_{\mathrm{max}}}$. We apply the sigmoid function $\sigma$ elementwise to obtain the probability that alternative $r$ is in the winning set. With $m$ the number of alternatives in profile $\boldsymbol{P}$, we define the decoding function
$d_{m}(\hat{y})$ :=$:=\big{\{}r\in\{0,\ldots,m\}:\sigma(\hat{y}_{r})>0.5\big{\}}.$
@@ 4.4 Loss Functions
For each axiom, we also design a loss function that enforces satisfaction of that axiom. Concretely, for the anonymity axiom, this is done as follows. Given the network $f_{w}$ and profile $\boldsymbol{P}$, uniformly sample $N$-many permutations $\pi_{1},\ldots,\pi_{N}$ of the set of voters of $\boldsymbol{P}$ and define
@@
where $\mathrm{KL}$ is Kullback–Leibler divergence. For the other axioms, we proceed similarly (see Appendix A, where we also discuss differentiability).
4.5 Evaluation Metrics
We have two ways of evaluating the model: accuracy and axioms. First, we calculate the accuracy of the trained neural network on a given test set in two ways: Identity (or hard) accuracy is the percentage of pairs $(\boldsymbol{P},S)$ in the test set for which $F_{w}(\boldsymbol{P})=S$. Subset (or soft) accuracy is defined in the same way but replacing the identity with $F_{w}(\boldsymbol{P})\subseteq S$. Second, we calculate the satisfaction degrees for the various axioms of the voting rule $F_{w}$ that the trained neural network realizes (see Section 5.3 for the details).
5 Experimental Setup
5.1 Voting-Theoretic Parameters
We work with $n_{\mathrm{max}}=77$ and $m_{\mathrm{max}}=7$ and all four profile distributions in the first experiment and with $n_{\mathrm{max}}=55$ and $m_{\mathrm{max}}=5$ and with IC and Mallows in the other experiments. The first experiment does not show a qualitative difference between these settings, but the latter is computationally more efficient.
We use the Mallows distribution with a parameter rel-$\phi$ (randomly generated) that, together with the number of alternatives, determines the value of the dispersion parameter $\phi$. According to Boehmer et al., (2021) and Boehmer et al., (2023), this methodology generates data that resemble more closely those of real elections. We use the Urn-R distribution (Boehmer et al.,, 2021), where, for each generated profile, $\alpha$ is chosen according to a Gamma distribution with shape parameter $k=0.8$ and scale parameter $\theta=1$. The other distributions do not need further parameters.
5.2 Synthetic Data Generation
@@
5.3 Evaluating Axiom Satisfaction
@@
To evaluate the axiom satisfaction of a voting rule (be it realized by a neural network or an existing one), we sample $400$ profiles on which the axioms are applicable. We use the same profile distribution $\mu$ as was used for training the neural network, and we again randomly choose integers $n\in[1,n_{\mathrm{max}}]$ and $m\in[1,m_{\mathrm{max}}]$ before $\mu$-sampling a profile with $n$ voters and $m$ alternatives. To compute whether an axiom is satisfied for a profile, the axioms of anonymity, neutrality, and independence require sampling of permutations. We sample, per profile, $50$, $50$, and $4^{n}$ (with $n$ the number of voters in the considered profile) permutations, respectively.555 Independence requires both a permutation of voters and of alternatives, hence we sample more permuted versions of the given profile.
5.4 Hyperparameters
@@
All models use ReLU as activation function. Our MLP has three hidden layers with 128 neurons each. The CNN has two convolution layers with kernel size $(5,1)$ and $(1,5)$, respectively (and 32 or 64 channels), followed by three linear layers with 128 neurons. The WEC has the word embedding layer, then the averaging layer, and then three linear layers with 128 neurons. For pre-training the embedding layer with word2vec, we use a corpus size of $10^{5}$, an embedding dimension of $200$, and a window size of $7$. [^6] That is in the setting $n_{\mathrm{max}}=77$ and $m_{\mathrm{max}}=7$. When $n_{\mathrm{max}}=55$ and $m_{\mathrm{max}}=5$, we reduce this to a corpus size of $2\times 10^{4}$, an embedding dimension of $100$, and a window size of $5$. This results in the following numbers of parameters in the setting $n_{\mathrm{max}}=77$ and $m_{\mathrm{max}}=7$: the MLP has $500.487$ parameters, the CNN has $1.834.439$ parameters, and the WEC has $1.226.143$ parameters. Thus, the models are roughly comparable in size.
@@
For training, we use the AdamW algorithm (Loshchilov and Hutter,, 2019). We use a batch size of 200. Since we have synthetic data, we do not use epochs and hence only specify the number of gradient steps. In experiment 1, 2, and 3, these are 15.000, 5.000, and 15.000, respectively. Similar to Anil and Bao, (2021), we use as a learning rate scheduler cosine annealing with warm restarts (Loshchilov and Hutter,, 2017). All results are reported for one fixed seed. (In the Appendix, tables 4 and 5 report averaged results across different seeds.) All experiments have been run on a laptop without GPU.
6 Results and Analysis

### v2

@@ 3.2. Axioms
We define axioms as functions that map a voting rule and a preference profile to a value in $\{-1,1,0\}$, where 0 means that the axiom is not applicable, 1-1 means that the axiom is violated, and 11 that it is satisfied. The satisfaction degree of a rule with respect to a given axiom is the ratio of the number of sampled profiles in which the axiom is satisfied to the number of sampled profiles in which it is applicable. We focus on axioms that capture basic and diverse normative properties of a voting rule $F$.
@@
All voting rules defined above satisfy anonymity and neutrality, as well as the Pareto principle, for all preference profiles. They all violate independence for some preference profile. Several of them such as Copeland, Llull, Blacks, Banks, Stable voting, and Weak Nanson always satisfy the Condorcet principle.
3.3. Distributions of Preference Profiles
Specifying the distribution of preference data is essential to studying the voting behavior of a society. Indeed, there is increasing interest within the computational social choice community towards a line of work called ‘map of elections’ that attempts to systematize simulation experiments relying on synthetic voting data (Boehmer et al., 2021, 2023). In this vein, we aim to ensure that our results are independent of the specific choice of the distribution. We thus employ four representative distributions that aim to cover elections of different nature, as explained below. In Section 5.1 we specify the choice of parameters of these distributions for our experiments, which allows them to capture a wide range of realistic scenarios. It is also important to note that all our experimental results are found to be robust across distributions; thus we do not expect that by examining additional distributions we would discover significantly different insights.
@@ 3.3. Distributions of Preference Profiles
Impartial Culture (IC) assumes that all preference profiles have the same probability of appearing. Each preference of a voter in a profile is sampled uniformly at random. The Mallows distribution (Mallows, 1957) fixes a reference ranking $P$ and assumes that each voter’s preference is close to that ranking. Closeness to the reference ranking is defined using the Kendall Tau distance, parameterized by a dispersion parameter $\phi\in(0,1]$.888 The Kendall Tau distance between two rankings $P$ and $Q$ over the same set of alternatives is the number of pairs of alternatives $(a,b)$ such that $a$ is preferred over $b$ in $P$ but not so in $Q$. This distribution reduces to IC when $\phi=1$ and concentrates all mass on $P$ as $\phi$ tends to 0.
@@
The IC and Mallows distributions are complementary: IC is simplistic and widely employed in theoretical works on voting rules as discussed earlier in the literature review; it captures an extreme case with no correlation between preferences of voters. Mallows is often employed in numerical studies of voting rules that use artificial data but wish to capture more realistic voting scenarios (Caragiannis and Micha, 2017; Lee et al., 2014).
@@
The next two distributions also capture more intricate relationships between the preferences in a profile. According to the 2D-Euclidean distribution, voters and alternatives are distributed randomly in 2-dimensional Euclidean space, and the closer an alternative is to a voter the more the voter prefers that alternative. Finally, the Urn distribution (Eggenberger and Pólya, 1923) generates a profile given a parameter $\alpha\in[0,\infty)$. Voters randomly draw their ranking from an urn. Initially, the urn includes all possible rankings over the alternatives. After a voter randomly draws from the urn, we add to the urn $\alpha n!$ copies of that ranking. When $\alpha=0$, this reduces to IC.
4. Method
@@
To answer our research questions, we develop the axiomatic deep voting framework, visualized in Figure 3. It is built around a neural network, which is a function $f_{w}:\mathbb{R}^{i}\to\mathbb{R}^{j}$ parametrized by weights $w\in\mathbb{R}^{k}$. We will instantiate this with three different neural network architectures (see Section 4.1). Every profile $\boldsymbol{P}$ is mapped, via an encoding function $e$ (see Section 4.2), to a vector $x=e(\boldsymbol{P})\in\mathbb{R}^{i}$, for which the neural network produces an output $\hat{y}\in\mathbb{R}^{j}$. [^9] For our third architecture, the encoding function is part of the neural network, while for the first two it is independent (see Section 4.2); hence we treat $e$ as a separate entity here. The decoding function $d$ (see Section 4.3) turns this output into a winning set S=d(y^)S=d($\hat${y}). Thus, this setup realizes the voting rule:
@@
[…]
The axiomatic deep voting architecture in diagrammatic form.
Figure 3. The axiomatic deep voting architecture.
4.1. Architectures
@@
First, multi-layer perceptrons (MLPs)—also known as feed-forward neural network—are the classic deep neural net (see, e.g., Goodfellow et al., 2016, ch. 6). They consist of an input layer of neurons, one or more hidden layers, and an output layer.
Second, convolutional neural networks (CNNs) are a standard architecture to process grid-like input data such as images (see, e.g., Goodfellow et al., 2016, ch. 9), and in our case profiles. Compared to MLPs, they additionally use so-called convolutional layers to capture local, invariant patterns in the input.
Third, we devise an architecture that satisfies the anonymity axiom by design: We view profiles as sentences whose words are the rankings. We use the word embedding algorithm Word2vec (Mikolov et al., 2013) to map each ranking to a high-dimensional embedding vector. These vectors are averaged—hence we get anonymity—and an MLP then classifies this average into a winning set. This combined architecture we call here word embedding classifiers (WECs).
4.2. Encoding
@@
How should we encode $\tilde{\boldsymbol{P}}$ so it can be inputted to a neural network? The most straightforward way is to read each alternative $a_{r}^{s}\in M$ as the number that it is and the padding symbol $\sim$ as, say, 1-1. Then the matrix $\tilde{\boldsymbol{P}}$ is regarded as a vector of dimension $m_{\mathrm{max}}n_{\mathrm{max}}$. However, this does not perform well, so, following Anil and Bao (2021), we represent an alternative not as a number but as a one-hot vector. For $a\in\{0,\ldots,m_{\mathrm{max}}-1\}$, let $\overline{a}$ be the vector of length $m_{\mathrm{max}}$ that is $1$ at position $a$ and 0 everywhere else. For the padding symbol, let $\overline{\sim}$ be the vector of length $m_{\mathrm{max}}$ that is 0 everywhere. We write $\overline{\boldsymbol{P}}=(\overline{\tilde{a}_{r}^{s}})_{r,s}$.
@@
The encoding function for CNNs regards the matrix $\overline{\boldsymbol{P}}$ as a pixel image: the ‘pixel’ at position $(r,s)$ has the ‘color value’ $\overline{\tilde{a}_{r}^{s}}$. Thus, $e_{\mathrm{CNN}}$ maps profile $\boldsymbol{P}$ to the matrix $\overline{\boldsymbol{P}}$ recast as a tensor with dimensions $(\mathit{channel},\mathit{height},\mathit{width})=(m_{\mathrm{max}},m_{\mathrm{max}},n_{\mathrm{max}})$. This tensor can then be inputted into the CNN.
The encoding function for WECs regards the profile $\boldsymbol{P}=(P_{1},\ldots,P_{n})$ as a sentence with words $P_{i}$. We train it to embed these words into vectors of a fixed high dimension. Thus, unlike the previous encoding functions, this one is not separate from the neural network but rather forms the first layer of the WEC, with the remaining layers processing the embedding vectors. More precisely, we first pre-train the embeddings as follows. For a given corpus size $c$, we sample $c$-many profiles from a given distribution of profiles (e.g., IC) to form our corpus (i.e., a set of sentences). The rankings occurring in the profiles form the vocabulary of this corpus, to which we add the unk token (to later represent unknown rankings, i.e., rankings that are not in the vocabulary) and the pad token (to pad a profile to length $n_{\mathrm{max}}$). Due to the unk token, this encoding applies to all profiles, even if it contains rankings that are not part of the model’s vocabulary. Using Word2vec, we train embeddings which represent words in the vocabulary as vectors. When instantiating the WEC architecture, these embeddings form the first layer: it maps the profile $(P_{1},\ldots,P_{n})$ to the corresponding embedding vectors $(v_{1},\ldots,v_{n})$. The next layer averages these vectors into a single vector $v$, followed by several linear layers ending with the output layer.
4.3. Decoding
Given a profile $\boldsymbol{P}$ as input, all neural network architectures produce as output the logits $\hat{y}=(\hat{y}_{0},\ldots,\hat{y}_{m_{\mathrm{max}}})$ in $\mathbb{R}^{m_{\mathrm{max}}}$. We apply the sigmoid function $\mathsf{sig}$ elementwise to obtain the probability that alternative $r$ is in the winning set.101010 We use $\mathsf{sig}$’ instead of $\sigma$’ to denote the sigmoid function, in order to not confuse it with the previous use of $\sigma$’ for permutations of alternatives. With $m$ the number of alternatives in profile $\boldsymbol{P}$, we define the decoding function
@@ 4.4. Loss Functions
A main contribution of this paper is that, for each axiom, we also design a loss function that enforces satisfaction of that axiom. So, for each axiom $\mathsf{ax}$, we define a function $L_{\mathsf{ax}}(f_{w},\boldsymbol{P})$ that takes as input the function $f_{w}$ computed by the neural network with weights $w$ and a profile $\boldsymbol{P}$. It outputs a non-negative real number describing numerically how much the axiom is satisfied: 0 means perfect axiom satisfaction, while higher numbers mean worse axiom satisfaction. We now define these loss functions, which we will use later.
Anonymity. Given the network $f_{w}$ and profile $\boldsymbol{P}$, uniformly sample $N$-many permutations $\pi_{1},\ldots,\pi_{N}$ of the set of voters of $\boldsymbol{P}$ and define
@@
where $\mathrm{KL}$ is Kullback–Leibler divergence. [^12] Though, in principle, other distance/similarity functions can be considered.
Condorcet. If $\boldsymbol{P}$ has no Condorcet winner, $L_{C}(f_{w},\boldsymbol{P}):=0$, and otherwise, if that Condorcet winner is alternative $a$, define (recall $\overline{a}$ is the one-hot vector for alternative $a$)
Pareto. We define (recall that $n_{a\succ b}^{\boldsymbol{P}}=n$ means that all voters in $\boldsymbol{P}$ rank $a$ above $b$)
$L_{P}(f_{w},\boldsymbol{P}):=\sum_{a,b\text{ with }n_{a\succ b}^{\boldsymbol{P}}=n}\mathsf{sig}\Big{(}f_{w}(e(\boldsymbol{P}))_{b}\Big{)}.$
Independence. Define $L_{I}(f_{w},\boldsymbol{P}):=0$ if $\boldsymbol{P}$ does not have at least two alternatives. Otherwise, randomly sample $N$-many pairs $(a_{r},b_{r})$ of distinct alternatives in $\boldsymbol{P}$ and randomly sample, for each ranking $P_{k}$ of $\boldsymbol{P}=(P_{1},\ldots,P_{n})$, a shuffling $P^{\prime}_{k}$ of $P_{k}$ in which, however, the order of $a_{r}$ and $b_{r}$ is the same as in $P_{k}$, and set $\boldsymbol{P}_{r}:=(P^{\prime}_{1},\ldots,P^{\prime}_{n})$. Write $\hat{y}:=f_{w}(e(\boldsymbol{P}))$ and $\hat{y}^{r}:=f_{w}(e(\boldsymbol{P}_{r}))$, and define
No winner. Recall that voting rules are required to output at least one winner. This is usually not called an axiom, and we did not hard-code this into our architectures. So we also want to optimize our neural networks to align with this requirement. Hence we define the ‘no winner’ loss as follows. Writing $\hat{y}=f_{w}(e(\boldsymbol{P}))$, we want that at least one of the numbers in $p:=\big{(}\mathsf{sig}(\hat{y}_{1}),\ldots,\sigma(\hat{y}_{m})\big{)}$ is above $0.5$, i.e., the maximum norm $\|p\|_{\infty}$ should be above $0.5$. Hence the more it is below that, the worse the loss:
$L_{NW}(f_{w},\boldsymbol{P}):=\max\big{(}0.5-\|p\|_{\infty},0\big{)}.$
4.5. Evaluation Metrics
We have two ways of evaluating the model: accuracy and axioms. First, we calculate the accuracy of the trained neural network on a given test set in two ways: Identity (or hard) accuracy is the percentage of pairs $(\boldsymbol{P},S)$ in the test set for which $F_{w}(\boldsymbol{P})=S$. Subset (or soft) accuracy is defined in the same way but replacing the identity with $F_{w}(\boldsymbol{P})\subseteq S$. Second, we calculate the satisfaction degrees for the various axioms of the voting rule $F_{w}$ that the trained neural network realizes (see Section 5.3 for details).
5. Experimental Setup
We describe all details for designing and evaluating our experiments.
5.1. Voting-Theoretic Parameters
We work with all four profile distributions and with $n_{\mathrm{max}}=77$ and $m_{\mathrm{max}}=7$ in the first experiment and $n_{\mathrm{max}}=55$ and $m_{\mathrm{max}}=5$ in the other experiments. The first experiment does not show a qualitative difference between these settings, but the latter is computationally more efficient.
For all experiments, we use the Mallows distribution with a parameter rel-$\phi$ (randomly generated) that, together with the number of alternatives, determines the value of the dispersion parameter $\phi$. According to (Boehmer et al., 2021) and (Boehmer et al., 2023), this methodology generates data that more closely resemble those of real elections. We use the Urn-R distribution (Boehmer et al., 2021), where, for each generated profile, $\alpha$ is chosen according to a Gamma distribution with shape parameter $k=0.8$ and scale parameter $\theta=1$. The other distributions do not need further parameters.
5.2. Synthetic Data Generation
@@
5.3. Evaluating Axiom Satisfaction
@@
To evaluate the axiom satisfaction of a voting rule (be it realized by a neural network or an existing one), we sample $400$ profiles on which the axioms are applicable. We use the same profile distribution $\mu$ as was used for training the neural network, and we again randomly choose integers $n\in[1,n_{\mathrm{max}}]$ and $m\in[1,m_{\mathrm{max}}]$ before $\mu$-sampling a profile with $n$ voters and $m$ alternatives. To compute whether an axiom is satisfied for a profile, the axioms of anonymity, neutrality, and independence require sampling of permutations. We sample, per profile, $50$, $50$, and w(mw)256w(m-w)256 permutations, respectively (where $w$ is the number of winners according to the rule on the profile, and hence mwm-w is the number of losers).141414 Independence considers more ‘degrees of freedom’, so we take more samples. Specifically, we go through all pairs $(x,y)$ where $x$ is a winner and $y$ a loser, and we sample, for each voter, $4$ alternative rankings that, however, have the same relative order of $x$ and $y$ as the actual ranking submitted by that voter, and then build $4^{4}=256$ profiles out of these alternative rankings and check that $y$ still does not win.
5.4. Hyperparameters
@@
All models use ReLU as the activation function. Our MLP has four hidden layers with 128 neurons each, like those of (Anil and Bao, 2021). The CNN has two convolution layers with kernel size $(5,1)$ and $(1,5)$, respectively (and 32 or 64 channels), followed by three linear layers with 128 neurons. Thus, the first kernel can pick up local patterns in the rankings of the voters, while the second kernel can pick up local patterns among the $i$-th preferred alternatives of the voters. (Appendix A.2 establishes the optimality of this choice when compared to other kernel sizes and additional pre-processing.) The WEC has the word embedding layer, then the averaging layer, and then three linear layers with 128 neurons. For pre-training the word embedding layer with word2vec, we use a corpus size of $10^{5}$, an embedding dimension of $200$, and a window size of 77.151515That is in the setting $n_{\mathrm{max}}=77$ and $m_{\mathrm{max}}=7$. When $n_{\mathrm{max}}=55$ and $m_{\mathrm{max}}=5$, we reduce this to a corpus size of $2\times 10^{4}$, an embedding dimension of $100$, and a window size of $5$. The corpus size is chosen large enough so that no occurrences of the unk token are observed in $1,000$ sampled profiles.
This results in the following numbers of parameters in the setting $n_{\mathrm{max}}=77$ and $m_{\mathrm{max}}=7$: $500,487$ (MLP), $1,834,439$ (CNN), and $1,226,143$ (WEC). In the setting $n_{\mathrm{max}}=55$ and $m_{\mathrm{max}}=5$ this reduces to: $193,285$ (MLP), $232,165$ (CNN), and $45,585$ (WEC). Thus, the models have roughly comparable capacities. Section A in the Appendix motivates these choices via hyperparameter tuning.
@@
For training, we use the AdamW algorithm (Loshchilov and Hutter, 2019). We use a batch size of 200. Since we have synthetic data, we do not use epochs and hence only specify the number of gradient steps. In experiments 1, 2, and 3, these are 15,000, 5,000, and 15,000, respectively. Similar to Anil and Bao (2021), we use as a learning rate scheduler cosine annealing with warm restarts (Loshchilov and Hutter, 2017). All results are reported for one fixed seed. (In the Appendix, Table 3 performs cross validation and Tables 5 and 8 report averaged results across different seeds.) All experiments were run on a laptop without GPU.
6. Results and Analysis